OpenStudy (anonymous):

use the elimination method to solve the system -4x+4y=-8 x-4y=-7 options A-(3,5) B-(5,3) C-(-3,-5) D-(-5,-3) Please show work

OpenStudy (anonymous):

-4x+4y=-8 x-4y=-7 ---------- -3x=-15 x=5

OpenStudy (anonymous):

how do u get the second part

OpenStudy (anonymous):

what second part?

OpenStudy (anonymous):

the second answer

OpenStudy (anonymous):

you mean y-coordinate?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

I dont know whats the name of this method in english but thats what i use to solve: x=-7+4y -4(-7+4y)+4y=-8 28-16y+4y=-8 -12y=-8-28 y=-36/-12 y=3

OpenStudy (anonymous):

(5,3)

OpenStudy (anonymous):

x-4y=-7 as x=5, 5-4y=-7 -4y=-7-5 -4y=-12 y=3

OpenStudy (anonymous):

thanks so much that was very helpful

OpenStudy (anonymous):

-4x+4y = -8........(i) x-4y = -7...........(ii) add equation (i) and (ii) -4x+x+4y-4y = -8-7 -3x = -15 x = 5 substituting x= 5 in (ii) 5+7 =4y y = 12/4 = 3 (5,3)

OpenStudy (anonymous):

@matias.mingo your method is substitution method as you substitute one variable in place of other.

OpenStudy (anonymous):

you are correct sir